If one number is 4 times another, and their product is 4 times their sum, what are the two numbers needed for the solution? I came up with this formula, but something is missing: x=one number and 4x + x = the other
4x+x = 4(4x+x)
5x = 20x

Answer by ewatrrr(24785)


Hi
Letting x and 4x represent the numbers..good work.
Writing as we Read, important to get the words into a written equation***
product = 4 times their sum
x(4x) = 4(x + 4x) |may be tendency to over think w/o writing while reading
4x^2 -20x = 0
4x(x-5) = 0 x = 0 and x = 5 |tossing out x = 0
x = 5, the other numbers is 20 (4*5)
CHECKING our Answer***
100 = 4*25